# Multi-Robot Object SLAM using Distributed Variational Inference

### v1

@@ V-E Feature initialization
In this way, we can augment the current state and covariance matrix as follows:
@@ VI Evaluation
Figure 2: Visualization of the trajectory and object estimates obtained by 3 robots on KITTI sequences 00, 05, 06 and 08.
We implement our distributed MSCKF algorithm with only stereo visual inputs and evaluate it by conducting experiments using the real-world KITTI dataset [40] and simulated data involving a larger number of robots.
@@ VI-A2 Geometric features
Same as [38], we extract geometric features using the FAST corner detector [5]. The KLT optical flow algorithm [41] is applied to track the features temporally and match them across stereo images. Outlier rejection is performed using 2-point RANSAC for temporal tracking and the known essential matrix for stereo matching. Finally, circular matching similar to [42] is performed to further remove outliers.
@@ VI-A5 Results and analysis
Qualitative results from three-robot collaborative object SLAM on the KITTI dataset are shown in Fig. 2. We show the root mean square error (RMSE) of the robot trajectory estimates in Table I and the mean distances between estimated object positions and the ground truth in Table II. We do use alignment for the trajectory RMSE [44] because trajectory transformations affect the object mapping errors. We can see that, although consensus averaging can harm the estimation accuracy for some robots compared to running individual MSCKF algorithms at each robot, it helps improve the overall team performance in terms of both localization and object mapping. Intuitively, the performance of the best robot in the team is sacrificed to improve the overall team performance in localization and mapping. The distributed MSCKF achieves better agreement in the map estimates among the different robots. Although it might be obvious that the map differences reduce with consensus averaging, we still compare the average object position differences with and without averaging in Table III to quantify the reduction in disagreement. We also claim that the consensus averaging step does not add much time overhead because the robots communicate only common landmarks, meaning that the corresponding covariance $\Sigma^{\mathbf{y}}_{i,t}$ in (19) is small, and only perform the averaging with the one-hop communication neighbors. The computation time used by different components in the algorithm is shown in Fig. 3. We can see that consensus averaging takes a small portion of time compared with the MSCKF update.
@@ VI-B Simulated data
To test the performance of our algorithm with increasing numbers of robots, we generated simulated data for 3-15 robots.
@@ VI-B2 Results
Figure 4: Visualization of the trajectory and object estimates obtained by 15 robots in simulation.
A visualization of the simulation result is provided in Fig. 4. The quantitative results from the simulations with fully connected graph are shown in Table IV. We can see that our algorithm scales efficiently with an increasing number of robots in the team while continuing to outperform decoupled MSCKF algorithms run at each robot individually. We also analyze the effect of connectivity in Fig. 5. In the experiment, each robot loses communication with each neighbor according to a certain rate. We can see that our algorithm is robust to communication lost rate up to $0.9$. We see that with small numbers of robots (3 and 5), the errors oscillate as the lost rate increases but with relatively large numbers of robots (10 and 15), the errors seem to decrease as the lost rate increases. Analyzing the effect of randomly connected graphs e.g. broad gossip [45] can be future work.
Figure 5: Analysis of the effect of the robot network connectivity.
@@
Separate trajectory RMSE (m) 1.114 1.091 0.786 0.766
Consensus trajectory RMSE (m) 0.594 0.479 0.419 0.614
Separate object error (m) 1.131 1.275 1.034 0.975
Consensus object error (m) 0.524 0.470 0.482 0.589
@@ ACKNOWLEDGMENTS
The authors are grateful to Shubham Kumar, Shrey Kansal, and Kishore Nukala from the University of California San Diego for technical discussions and help with preparing the datasets for the evaluation.
@@ A-B Proof of Proposition 3
Note that $\frac{\delta c_{i}}{\delta q_{i}}(q_{i}^{(k)})$, $\log q_{i}^{(k)}$ and $\log q_{i}^{(k)}(\mathbf{y})$ are all functions of $\mathbf{s}_{i}$. We have $\frac{\delta c_{i}}{\delta q_{i}}=-\log p_{i}(\mathbf{s}_{i},\mathbf{z}_{i},\mathbf{u}_{i})+\log q_{i}(\mathbf{s}_{i})+1$ [47, Proposition 1] and $\log p_{i}(\mathbf{s}_{i},\mathbf{z}_{i},\mathbf{u}_{i})$ can be decomposed with (3) so that (30) can be rewritten as
@@
The result can also be interpreted by Proposition 1, (38) and (39) are geometric averaging of the marginal Gaussians over $\mathbf{y}$ [48, Lemma 1] which correspond to $\prod_{j\in\mathcal{N}_{i}}[q^{(k)}_{j}(\mathbf{y})]^{A_{ij}}$ in (9). (41) is constructing a new joint distribution with the averaged marginal (Proposition 2) which correspond to $q_{i}^{(k)}(\mathbf{x}_{i}|\mathbf{y})\prod_{j\in\mathcal{N}_{i}}[q^{(k)}_{j}(\mathbf{y})]^{A_{ij}}$ in (9). The step term $\alpha_{k}(\bar{F}^{-\top}_{i}\bar{W}_{i}^{-1}\bar{F}^{-1}_{i}\!+\bar{H}^{\top}_{i}\bar{V}_{i}^{-1}\bar{H}_{i}-\!\Sigma^{-1}_{i,(k)})$ and $\alpha_{k}(\bar{F}_{i}^{-\top}\bar{W}_{i}^{-1}\bar{G}_{i}\mathbf{u}_{i}+\bar{H}_{i}^{\top}\bar{V}_{i}^{-1}\mathbf{z}_{i}-\Sigma^{-1}_{i,(k)}\boldsymbol{\mu}_{i,(k)})$ correspond to $[p_{i}(\mathbf{x}_{i},\mathbf{y},\mathbf{z}_{i},\mathbf{u}_{i})/q_{i}^{(k)}(\mathbf{x}_{i},\mathbf{y})]^{\alpha_{k}}$ in (9).

### v2

@@ V-E Feature initialization
Thus, we can augment the current state and covariance:
@@
Figure 2: Trajectory and object estimates of (a) 3 robots on KITTI sequences 00, 05, 06 and 08, (b) 15 robots in simulation.
@@ VI Evaluation
We implemented the distributed MSCKF using only stereo camera observations and evaluated it on the KITTI dataset [41] and on a simulated dataset with a larger number of robots. All experiments were carried out on a laptop with i9-11980HK@2.60 CPU, 16 GB RAM, and RTX 3080 GPU.
@@ VI-A2 Geometric features
We extract geometric features using the FAST corner detector [5]. The KLT optical flow algorithm [42] is used to track the features across stereo images. Outlier rejection is performed using 2-point RANSAC for temporal tracking and the known essential matrix for stereo matching. Finally, circular matching similar to [43] is performed to further remove outliers.
@@ VI-A4 Odometry
Figure 4: Analysis of the effect of the robot network connectivity. The dashed lines show RMSE without averaging.
@@ VI-A5 Results and analysis
We found empirically that setting the correlation matrix (25) to zero gives the best results. We assume that this is because libviso2 [38] uses SURF features [45], while the update step is performed using FAST features [5] and the correlation is negligible. Qualitative results from three-robot collaborative object SLAM on the KITTI dataset are shown in Fig. 2(a). We show the root mean square error (RMSE) of the robot trajectory estimates in Table I and the mean distances between estimated object positions and the ground truth in Table II. We do not use alignment for the trajectory RMSE [46] because trajectory transformations affect the object mapping errors. Some ways to mitigate the effect of bad estimates include resilient consensus [47] or adaptive adjacency weights $A_{ij}$ depending on the robots’ measurement accuracy. Although consensus averaging can harm the estimation accuracy for some robots compared to running individual MSCKF algorithms for each robot, it helps improve the overall team performance in both localization and object mapping. The separate MSCKF sometimes perform worse than libviso2 [38]. This is because the object observation is too noisy. Updating with only object features can give an error up to 10 times as in Table I. The distributed MSCKF achieves better agreement in the map estimates among the robots. We compare the object position differences with and without averaging in Table III to quantify the reduction in disagreement. We also claim that the consensus averaging step does not add much time overhead because the robots communicate only common landmarks, meaning that the corresponding covariance $\Sigma^{\mathbf{y}}_{i,t}$ in (19) is small, and only perform averaging with one-hop neighbors. The computation time used by different components in the algorithm is shown in Fig. 3. Consensus averaging takes a small portion of time compared with the MSCKF update.
@@ VI-B Simulated data
To test our algorithm with increasing numbers of robots, we generated simulated data for 3-15 robots.
@@ VI-B2 Results
The results are visualized in Fig. 2(b). The quantitative results from the simulations with a fully connected graph are shown in Table IV. Our algorithm scales efficiently with an increasing number of robots while continuing to outperform decoupled MSCKF algorithms for each robot. We also analyze the effect of connectivity in Fig. 4. In the experiment, each robot loses communication with each neighbor according to rate $r$, i.e., each edge in the graph at each time step is removed independently with probability $r$. We see that our algorithm is robust to communication loss rate up to $r=0.9$. With small numbers of robots (3 and 5), the errors oscillate as the loss rate increases but with relatively large numbers of robots (10 and 15), the errors seem to decrease as the loss rate increases. Analyzing the effect of randomly connected graphs, e.g., broad gossip [48], will be considered in future work.
@@
Separate trajectory RMSE (m) 1.318 1.452 1.464 1.385
Consensus trajectory RMSE (m) 0.605 0.748 0.941 0.933
Separate object error (m) 1.402 1.604 1.605 1.573
Consensus object error (m) 0.514 0.536 0.752 0.737
@@ ACKNOWLEDGMENTS
The authors are grateful to Shubham Kumar, Shrey Kansal, and Kishore Nukala from University of California San Diego for technical discussions and help with dataset preparation.
@@ A-B Proof of Proposition 3
Note that $\frac{\delta c_{i}}{\delta q_{i}}(q_{i}^{(k)})$, $\log q_{i}^{(k)}$ and $\log q_{i}^{(k)}(\mathbf{y})$ are all functions of $\mathbf{s}_{i}$. We have $\frac{\delta c_{i}}{\delta q_{i}}=-\log p_{i}(\mathbf{s}_{i},\mathbf{z}_{i},\mathbf{u}_{i})+\log q_{i}(\mathbf{s}_{i})+1$ [50, Proposition 1] and $\log p_{i}(\mathbf{s}_{i},\mathbf{z}_{i},\mathbf{u}_{i})$ can be decomposed with (3) so that (32) can be rewritten as
@@
The result can also be interpreted by Proposition 1, (40) and (41) are geometric averaging of the marginal Gaussians over $\mathbf{y}$ [51, Lemma 1] which correspond to $\prod_{j\in\mathcal{N}_{i}}[q^{(k)}_{j}(\mathbf{y})]^{A_{ij}}$ in (9). (43) is constructing a new joint distribution with the averaged marginal (Proposition 2) which correspond to $q_{i}^{(k)}(\mathbf{x}_{i}|\mathbf{y})\prod_{j\in\mathcal{N}_{i}}[q^{(k)}_{j}(\mathbf{y})]^{A_{ij}}$ in (9). The step term $\alpha_{k}(\bar{F}^{-\top}_{i}\bar{W}_{i}^{-1}\bar{F}^{-1}_{i}\!+\bar{H}^{\top}_{i}\bar{V}_{i}^{-1}\bar{H}_{i}-\!\Sigma^{-1}_{i,(k)})$ and $\alpha_{k}(\bar{F}_{i}^{-\top}\bar{W}_{i}^{-1}\bar{G}_{i}\mathbf{u}_{i}+\bar{H}_{i}^{\top}\bar{V}_{i}^{-1}\mathbf{z}_{i}-\Sigma^{-1}_{i,(k)}\boldsymbol{\mu}_{i,(k)})$ correspond to $[p_{i}(\mathbf{x}_{i},\mathbf{y},\mathbf{z}_{i},\mathbf{u}_{i})/q_{i}^{(k)}(\mathbf{x}_{i},\mathbf{y})]^{\alpha_{k}}$ in (9).
A-C Derivation of the Kalman filter with noise correlation
Suppose that there is correlation between the motion and observation noise:
$S=\mathbb{E}[\mathbf{w}_{t-1}\mathbf{v}_{t}^{\top}],$ (45)
and take two timesteps of only one robot’s state into account, namely, $\mathbf{s}_{t-1}$ and $\mathbf{s}_{t}$, by defining the following lifted terms
$\mathbf{s}\!=\![\mathbf{s}_{t-1}^{\top}\ \mathbf{s}^{\top}_{t}]^{\top}\!\!,\mathbf{u}^{\top}\!\!=\![\boldsymbol{\mu}_{t-1}^{\top}\ \mathbf{u}_{t-1}^{\top}]^{\top}\!\!,\bar{H}\!=\![0\ H],\bar{S}\!=\![0\ S^{\top}]^{\top}\!\!,$
$\bar{W}=\begin{bmatrix}\Sigma_{t-1}&0\\ 0&W\end{bmatrix},\bar{F}=\begin{bmatrix}I&0\\ F&I\\ \end{bmatrix},\bar{G}=\begin{bmatrix}I&0\\ 0&G\end{bmatrix},$ (46)
where we omit the subscript $i$ since there is only one robot. The term $\log p(\mathbf{s},\mathbf{z}_{t},\mathbf{u})$ in (5) can be re-written as:
$\log\!p(\mathbf{s},\mathbf{z}_{t},\mathbf{u})\!\!\propto\!-\frac{1}{2}\!\begin{bmatrix}\bar{F}^{-1}\mathbf{s}\!-\!\bar{G}\mathbf{u}\\ \mathbf{z}_{t}\!-\!\bar{H}\mathbf{s}\end{bmatrix}\!\!\begin{bmatrix}\bar{W}&\bar{S}\\ \bar{S}^{\top}&V\end{bmatrix}^{-1}\!\!\begin{bmatrix}\bar{F}^{-1}\mathbf{s}\!-\!\bar{G}\mathbf{u}\\ \mathbf{z}_{t}\!-\!\bar{H}\mathbf{s}\end{bmatrix}^{\top}\!\!\!\!.$ (47)
The derivatives in (32), i.e., $\frac{\partial g_{i}^{(k)}}{\partial\boldsymbol{\mu}^{\top}_{i}}$, $\frac{\partial g_{i}^{(k)}}{\partial\Sigma_{i}}$, and $\frac{\partial g_{i}^{(k)}}{\partial\Sigma_{i}}$ remain the same format except $c_{i}(q_{i})$ in (5) changes. Therefore, we can follow the exact derivation steps in Sec. A-B, which we omit and directly show that (17) becomes
$\bar{\Sigma}^{-1}$ $=\bar{F}^{-\top}A^{-1}\bar{F}+\bar{H}^{\top}C^{\top}\bar{F}^{-1}+\bar{F}^{-\top}C\bar{H}+\bar{H}B^{-1}\bar{H},$
$\bar{\Sigma}^{-1}\bar{\boldsymbol{\mu}}$ $=\bar{F}^{-\top}\bar{A}^{-1}\bar{G}\bar{\mathbf{u}}+\bar{H}\bar{C}^{\top}\bar{\mathbf{u}}+\bar{F}^{-\top}\bar{C}\mathbf{z}_{t}+\bar{H}^{\top}\bar{B}^{-1}\mathbf{z}_{t},$ (48)
where
$\bar{A}=\bar{W}-\bar{S}V^{-1}\bar{S}^{\top},\ \bar{B}=V-\bar{S}^{\top}\bar{W}^{-1}\bar{S},\bar{C}=-\bar{A}^{-1}\bar{S}V^{-1}.$ (49)
By substituting (49) and (46) into (A-C), we get
$\bar{\Sigma}^{-1}=$
$\begin{bmatrix}\Sigma_{t-1}^{-1}\!+\!F^{\top}\!\!A^{-1}F&F^{\top}\!A^{-1}SV^{-1}H\!-\!F^{\top}\!\!A^{-1}\\ H^{\top}\!V^{-1}\!S^{\top}\!A^{-1}\!F\!-\!A^{-1}\!F&A^{-1}\!+\!H^{\top}\!\!B^{-1}H...\\ \ &-\!H^{\top}\!V^{-1}\!S^{\top}\!\!A^{-1}\!-\!A^{-1}SV^{-1}H\end{bmatrix}\!\!,$
$\bar{\Sigma}^{-1}\bar{\boldsymbol{\mu}}=$
$\begin{bmatrix}\Sigma_{t-1}^{-1}\boldsymbol{\mu}_{t-1}\!-\!F^{\top}A^{-1}G\mathbf{u}_{t}\!+\!F^{\top}A^{-1}SV^{-1}\mathbf{z}_{t}\\ A^{-1}G\mathbf{u}_{t}\!-\!H^{\top}V^{-1}S^{\top}A^{-1}G\mathbf{u}_{t}\!-\!A^{-1}SV^{-1}\mathbf{z}_{t}\!+\!H^{\top}B^{-1}\mathbf{z}_{t}\end{bmatrix}\!\!.$ (50)
where
$A=W-SV^{-1}S^{\top},\ B=V-S^{\top}W^{-1}S.$ (51)
Following the approach in Eq. 3.113 in [37, Chapter 3.3.2], we multiply both sides of (50) by the matrix:
$\begin{bmatrix}I&0\\ (A^{-1}F-H^{\top}V^{-1}S^{\top}A^{-1}F)(\Sigma_{t-1}^{-1}+F^{\top}A^{-1}F)^{-1}&I\end{bmatrix}$ (52)
to obtain:
$\quad\begin{bmatrix}\Sigma_{t-1}^{-1}\!+\!F^{\top}\!A^{-1}F&F^{\top}\!A^{-1}SV^{-1}H\!-\!F^{\top}\!A^{-1}\\ 0&\Sigma_{t}^{-1}\end{bmatrix}\begin{bmatrix}\hat{\boldsymbol{\mu}}_{t-1}\\ \boldsymbol{\mu}_{t}\end{bmatrix}$
$=\begin{bmatrix}\Sigma_{t-1}^{-1}\boldsymbol{\mu}_{t-1}-F^{\top}A^{-1}G\mathbf{u}_{t}+F^{\top}SV^{-1}\mathbf{z}_{t}\\ \Sigma_{t}^{-1}\boldsymbol{\mu}_{t}\end{bmatrix},$ (53)
where $\hat{\boldsymbol{\mu}}_{t-1}$ is different from the estimation $\boldsymbol{\mu}_{t-1}$ at time $t-1$ since it is estimated with future observation, i.e. $\mathbf{z}_{t}$ and
$\Sigma_{t}^{-1}\!=$ $\!-\!(H^{\top}V^{-1}S^{\top}A^{-1}F-A^{-1}F)(\Sigma_{t-1}^{-1}+F^{\top}A^{-1}F)^{-1}$
$\ (F^{\top}A^{-1}SV^{-1}H-F^{\top}A^{-1})+A^{-1}+H^{\top}B^{-1}H$
$\ -H^{\top}V^{-1}S^{\top}A^{-1}-A^{-1}SV^{-1}H,$ (54)
$\Sigma_{t}^{-1}\boldsymbol{\mu}_{t}\!$ $=\!(H^{\top}\!V^{-1}S^{\top}\!A^{-1}F-A^{-1}F)(\Sigma_{t-1}^{-1}\!+\!F^{\top}\!A^{-1}F)^{-1}$
$(\Sigma_{t-1}^{-1}\boldsymbol{\mu}_{t-1}\!-\!F^{\top}\!A^{-1}\!G\mathbf{u}_{t}\!+\!F^{\top}\!A^{-1}\!SV^{-1}\!\mathbf{z}_{t})\!+\!A^{-1}\!G\mathbf{u}_{t}$
$-H^{\top}V^{-1}S^{\top}A^{-1}G\mathbf{u}_{t}-A^{-1}SV^{-1}\!\mathbf{z}_{t}+H^{\top}B^{-1}\!\mathbf{z}_{t}.$ (55)
To proceed, we need to use the Woodbury identity [52]:
$(A\!+\!CBC^{\top})^{-1}\!=\!A^{-1}\!-\!A^{-1}\!C(B^{-1}\!+\!C^{\top}\!A^{-1}\!C)^{-1}\!C^{\top}\!A^{-1},$ (56)
$(P^{-1}\!+\!B^{\top}\!R^{-1}\!B)^{-1}B^{\top}\!R^{-1}\!=\!PB^{\top}(BPB^{\top}\!+\!R)^{-1},$ (57)
where $P$ and $R$ should be positive definite. Focusing on the covariance first, i.e. (54), and using (56), we have
$A^{-1}-A^{-1}F(\Sigma_{t-1}^{-1}+F^{\top}A^{-1}F)^{-1}F^{\top}A^{-1}$ (58)
$=(A+F\Sigma_{t-1}F^{\top})^{-1}$
$=(W\!-\!SV^{-1}\!S^{\top}\!+\!F\Sigma_{t-1}\!F^{\top})^{-1}\!=\!(\Sigma_{t}^{+}\!-\!SV^{-1}S^{\top})^{-1},$
where $\Sigma^{+}_{t}$ is the predicted covariance. To shorten the equations, we define the following identitie:
$D=(\Sigma^{+}_{t}-SV^{-1}S^{\top})^{-1},\ E=(V-S^{\top}\Sigma^{+}_{t}S)^{-1}.$ (59)
This way, we get
$\Sigma_{t}^{-1}=$ $D+H^{\top}V^{-1}S^{\top}DSV^{-1}H+H^{\top}V^{-1}H$
$-H^{\top}V^{-1}S^{\top}D-DSV^{-1}H.$ (60)
Using (56) and (57), we get
$E\!=\!V^{-1}\!+\!V^{-1}\!S^{\top}\!DSV^{-1}\!\!,\ D\!=\!\Sigma_{t}^{+-1}\!+\!\Sigma_{t}^{+-1}\!SES^{\top}\!\Sigma_{t}^{+-1}$ (61)
$V^{-1}S^{\top}D=-ES^{\top}\Sigma^{+-1}_{t},\ DSV^{-1}=-\Sigma_{t}^{+-1}SE,$ (62)
and (60) becomes
$\Sigma_{t}^{-1}=$ $D+H^{\top}EH+H^{\top}ES^{\top}\Sigma_{t}^{+-1}+\Sigma_{t}^{+-1}SEH$
$=$ $\Sigma_{t}^{+-1}+\Sigma_{t}^{+-1}SES^{\top}\Sigma_{t}^{+-1}+H^{\top}EH$
$+H^{\top}ES^{\top}\Sigma_{t}^{+-1}+\Sigma_{t}^{+-1}SEH$
$=$ $\Sigma_{t}^{+-1}+(H^{\top}+\Sigma^{+-1}_{t}S)E(H+S^{\top}\Sigma^{+-1}_{t}).$ (63)
Letting $J=H+S^{\top}\Sigma^{+-1}_{t}$ and taking the inverse of the above equation by using (56), we get
$\Sigma_{t}=$ $\Sigma^{+}_{t}-\Sigma^{+}_{t}J^{\top}(J\Sigma^{+}_{t}J^{\top}+E^{-1})^{-1}J\Sigma^{+}_{t}$
$=$ $\Sigma^{+}_{t}-(\Sigma^{+}_{t}H^{\top}+S)(H\Sigma^{+}_{t}H^{\top}+V+HS+S^{\top}H^{\top})^{-1}$
$(H\Sigma^{+}_{t}+S^{\top}).$ (64)
Defining the Kalman gain as
$K_{t}=(\Sigma^{+}_{t}H^{\top}+S)(H\Sigma^{+}_{t}H^{\top}+V+HS+S^{\top}H^{\top})^{-1},$ (65)
the above equation can be rewritten as
$\Sigma_{t}=(I-K_{t}H)\Sigma_{t}^{+}-K_{t}S^{\top},$ (66)
Then, we show that $\boldsymbol{\mu}_{t}$ in (55) is equivalent to the following:
$\boldsymbol{\mu}_{t}=\boldsymbol{\mu}_{t}^{+}+K_{t}(\mathbf{z}_{t}-H\boldsymbol{\mu}^{+}_{t}).$ (67)
(57) tells us that
$A^{-1}F(\Sigma_{t-1}^{-1}+F^{\top}A^{-1}F)^{-1}\Sigma_{t-1}^{-1}\boldsymbol{\mu}_{t}$
$=(A+F\Sigma_{t-1}F^{\top})^{-1}F\boldsymbol{\mu}_{t}$
$=(W-SV^{-1}S^{\top}+F\Sigma_{t-1}F^{\top})^{-1}F\boldsymbol{\mu}_{t}$
$=(\Sigma^{+}_{t}-SV^{-1}S^{\top})^{-1}F\boldsymbol{\mu}_{t}=DF\boldsymbol{\mu}_{t}.$ (68)
With (58) and (68), we can simplify (55) as
$\Sigma_{t}^{-1}\boldsymbol{\mu}_{t}$ $=(I-H^{\top}V^{-1}S^{\top})(\Sigma_{t}^{+}-SV^{-1}S^{\top})^{-1}$
$\quad\;(F\boldsymbol{\mu}_{t-1}+G\mathbf{u}_{t}-SV^{-1}\mathbf{z}_{t})+H^{\top}V^{-1}\mathbf{z}_{t}$
$=(I\!-\!H^{\top}\!V^{-1}\!S^{\top})\!D(\boldsymbol{\mu}^{+}_{t}\!-\!SV^{-1}\!\mathbf{z}_{t})\!+\!H^{\top}\!V^{-1}\mathbf{z}_{t}.$ (69)
Then, our goal is to show that $\boldsymbol{\mu}_{t}$ in (67) and (69) are the same, i.e. the above equation can be written as:
$\Sigma_{t}^{-1}\boldsymbol{\mu}_{t}=\Sigma_{t}^{-1}\boldsymbol{\mu}_{t}^{+}+\Sigma_{t}^{-1}K_{t}(\mathbf{z}_{t}-H\boldsymbol{\mu}^{+}_{t})$ (70)
The Kalman gain can be rewritten by using (56) and (57) as
$K_{t}$ $=(\Sigma^{+}_{t}H^{\top}+S)(H\Sigma^{+}_{t}H^{\top}+V+HS+S^{\top}H^{\top})^{-1}$
$=\Sigma^{+}_{t}J^{\top}(J\Sigma_{t}^{+}J^{\top}+E^{-1})^{-1}$
$=(\Sigma_{t}^{+-1}+J^{\top}EJ)^{-1}J^{\top}E$
$=\Sigma_{t}J^{\top}E.$ (71)
Therefore,
$\Sigma^{-1}_{t}K_{t}(\mathbf{z}_{t}-H\boldsymbol{\mu}_{t})$
$=$ $J^{\top}E\mathbf{z}_{t}-J^{\top}EH\boldsymbol{\mu}_{t}$
$=$ $H^{\top}\!E\mathbf{z}_{t}\!+\!\Sigma_{t}^{+-1}\!SE\mathbf{z}_{t}\!-\!H^{\top}\!EH\!\boldsymbol{\mu}_{t}\!-\!\Sigma_{t}^{+-1}\!SEH\boldsymbol{\mu}_{t}.$ (72)
Substituting the first line of (63) into the right hand side of (70), we get
$\Sigma_{t}^{-1}\boldsymbol{\mu}_{t}^{+}+\Sigma_{t}^{-1}K_{t}(\mathbf{z}_{t}-H\boldsymbol{\mu}^{+}_{t})$
$=$ $D\boldsymbol{\mu}_{t}+H^{\top}EH\boldsymbol{\mu}_{t}+H^{\top}ES^{\top}\Sigma_{t}^{+-1}\boldsymbol{\mu}_{t}+\Sigma_{t}^{+-1}SEH\boldsymbol{\mu}_{t}$
$+H^{\top}\!E\mathbf{z}_{t}\!+\!\Sigma_{t}^{+-1}\!SE\mathbf{z}_{t}\!-\!H^{\top}\!EH\!\boldsymbol{\mu}_{t}\!-\!\Sigma_{t}^{+-1}\!SEH\boldsymbol{\mu}_{t}$
$=$ $D\boldsymbol{\mu}_{t}+H^{\top}ES^{\top}\Sigma_{t}^{+-1}\boldsymbol{\mu}_{t}+H^{\top}\!E\mathbf{z}_{t}+\Sigma_{t}^{+-1}SE\mathbf{z}_{t}$
$=$ $D\boldsymbol{\mu}_{t}-H^{\top}V^{-1}S^{\top}D\boldsymbol{\mu}_{t}$
$+H^{\top}V^{-1}\mathbf{z}_{t}+H^{\top}V^{-1}S^{\top}DSV^{-1}\mathbf{z}_{t}-DSV^{-1}\mathbf{z}_{t}$
$=$ $(I\!-\!H^{\top}\!V^{-1}\!S^{\top})D(\boldsymbol{\mu}^{+}_{t}\!-\!SV^{-1}\!\mathbf{z}_{t})+H^{\top}\!V^{-1}\mathbf{z}_{t},$ (73)
which aligns with (55), thus completing the proof.
In summary, given estimation from the previous timestep $\boldsymbol{\mu}_{t-1},\Sigma_{t-1}$, control input $\mathbf{u}_{t}$, and observation $\mathbf{z}_{t}$ at time $t$, the Kalman filter with correlated motion and observation noise is shown below.
Prediction:
$\boldsymbol{\mu}_{t}^{+}$ $=F\boldsymbol{\mu}_{t-1}+G\mathbf{u}_{t},$ (74)
$\Sigma_{t}^{+}$ $=F\Sigma_{t-1}F^{\top}+W.$ (75)
Update:
$K_{t}$ $=(\Sigma^{+}_{t}H^{\top}\!+\!S)(H\Sigma^{+}_{t}H^{\top}\!+\!V\!+\!HS\!+\!S^{\top}H^{\top})^{-1},$ (76)
$\boldsymbol{\mu}_{t}$ $=\boldsymbol{\mu}_{t}^{+}+K_{t}(\mathbf{z}_{t}-H\boldsymbol{\mu}_{t}^{+}),$ (77)
$\Sigma_{t}$ $=(I-\bar{K}_{t}H)\Sigma_{t}^{+}-K_{t}S^{\top},$ (78)
which aligns with [53, Table 5.1].